OpenStudy (anonymous):

(4x)^-2 positive exponent I came up with 1/62x^2

OpenStudy (anonymous):

\[(4x)^{-2}\] is the same as \[\frac{1}{(4x)^2}\]

OpenStudy (anonymous):

using this when you have \[(4x)^2=4^2x^2\]

OpenStudy (anonymous):

does this make sense

OpenStudy (anonymous):

\[a^{-n}=\frac{1}{a^n}\]

OpenStudy (anonymous):

so the answer would be 1/a^n?

OpenStudy (phi):

outkast's very first post gives you the answer. you could also write it as \[\frac{1}{16x^2} \]

OpenStudy (anonymous):

1/16x^2 can be wrote as a positive exponent?

OpenStudy (anonymous):

I guess your question is: Can (4x)^-2 be writen as a positive exponent? And the answer is: yes 1/(4x)^2 or 1/16x^2

OpenStudy (phi):

Not sure I understand your question: 1/16x^2 can be wrote as a positive exponent? It is a positive exponent. That is the answer to your original question.

OpenStudy (anonymous):

oh okay thanks. so I had the wrong answer 1/62x^2?

OpenStudy (phi):

the 62 or is that 6^2? is wrong. the 1/x^2 part is correct. you change \( (4x)^{-2} \) to \( \frac{1}{(4x)^{2}} \) the idea is not too complicated: flip the fraction and negate the exponent.

OpenStudy (phi):

so you can do \[ \frac{1}{(4x)^{-2} }= \frac{(4x)^2}{1} = (4x)^2 \] flip the fraction and negate the exponent. negate means change - to + or + to -